Leann just sold a $10,000 par value bond for $9,800. The bond interest rate was 7% per year payable quarterly. Leann owned the bond for 3 years. The 1st interest payment she received was 3 months after she bought the bond. She sold it immediately after receiving her 12th interest payment. Leann’s yield on the bond was 15% per year compounded quarterly. Determine the price she paid when she purchased the bond.
Answer #1

Bond's Selling Price = $9,800

Quarterly Coupon = (0.07/4)(10,000) = $175

Time Period = 12 quarters

Yield = 15%

Calculating Purchase Price of Bond,

Using TVM Calculation,

PV = [FV = 9,800, PMT = 175, N = 12, I = 0.15/4]

PV = $7,966.88

Bond Purchase Price = $7,966.88